15. Let





f(x,y) { 0, x² < y < 2x² 1. otherwise





Verify the existence of partial derivatives fr(0,0) and f(0, 0), and the differentiability of fat (0,0).

1
Expert's answer
2022-02-02T15:17:26-0500

From the given definition, we have that;

"\displaystyle
f(x,y)=\begin{cases}
0 &\text{if } x²<y<2x²\\
1&\text{if otherwise}
\end{cases}"

That is,

"\displaystyle
f(x,y)=\begin{cases}
0 &\text{if } y\in(x^2,2x^2)\\
1&\text{if }y\in(-\infty,x^2]\cup[2x^2,\infty)
\end{cases}"

Now, "\displaystyle
f(0,0)=1 \text{ since }0\in(-\infty,x^2]\cup[2x^2,\infty)"


Thus,


1] "\displaystyle
f_x(0,0)=\lim_{h\rightarrow0}\frac{f(h,0)-f(0,0)}{h}=\lim_{h\rightarrow0}\frac{f(h,0)-1}{h}"

But,"\displaystyle
\,\lim_{h\rightarrow0}\frac{f(h,0)-1}{h}\text{ exists iff }\lim_{h\rightarrow0^-}\frac{f(h,0)-1}{h}\text{ and }\lim_{h\rightarrow0^+}\frac{f(h,0)-1}{h}\text{ exists and are}\\
\text{equal. }"

Now, "\displaystyle
\lim_{h\rightarrow0^-}\frac{f(h,0)-1}{h}=\lim_{h\rightarrow0^-}\frac{1-1}{h}=0, \text{since }f(h,0) \rightarrow1\text{ as }h\rightarrow0^- \text{ because}\\"

"\displaystyle
0\in(-\infty,h^2]\cup[2h^2,\infty)"

Also, "\displaystyle
\lim_{h\rightarrow0^+}\frac{f(h,0)-1}{h}=\lim_{h\rightarrow0^+}\frac{1-1}{h}=0, \text{since }f(h,0) \rightarrow1\text{ as }h\rightarrow0^+ \text{ because}\\"

"\displaystyle
0\in(-\infty,h^2]\cup[2h^2,\infty)"

Since both left and right limit exists, we have;

"\displaystyle
f_x(0,0)=\lim_{h\rightarrow0}\frac{f(h,0)-f(0,0)}{h}=\lim_{h\rightarrow0}\frac{f(h,0)-1}{h}=0"


2] "\displaystyle
f_y(0,0)=\lim_{h\rightarrow0}\frac{f(0,h)-f(0,0)}{h}=\lim_{h\rightarrow0}\frac{f(0,h)-1}{h}"

But,

"\displaystyle
\,\lim_{h\rightarrow0}\frac{f(0,h)-1}{h}\text{ exists iff }\lim_{h\rightarrow0^-}\frac{f(0,h)-1}{h}\text{ and }\lim_{h\rightarrow0^+}\frac{f(0,h)-1}{h}\text{ exists and are}\\
\text{equal. }"

Now, "\displaystyle
\lim_{h\rightarrow0^-}\frac{f(0,h)-1}{h}=\lim_{h\rightarrow0^-}\frac{1-1}{h}=0, \text{since }f(0,h) \rightarrow1\text{ as }h\rightarrow0^- \text{ because}\\"

"\displaystyle
h\in(-\infty,0]\cup[0,\infty)"

Also, "\displaystyle
\lim_{h\rightarrow0^+}\frac{f(0,h)-1}{h}=\lim_{h\rightarrow0^+}\frac{1-1}{h}=0, \text{since }f(0,h) \rightarrow1\text{ as }h\rightarrow0^+ \text{ because}\\"

"\displaystyle
h\in(-\infty,0]\cup[0\infty)"

Since both left and right limit exists, we have;

"\displaystyle
f_y(0,0)=\lim_{h\rightarrow0}\frac{f(0,h)-f(0,0)}{h}=\lim_{h\rightarrow0}\frac{f(0,h)-1}{h}=0".


3] We need to check the continuity of the function at "\displaystyle
(0,0)." That is, we need to show that "\displaystyle
\lim_{(x,y)\rightarrow(0,0)}f(x,y)=f(0,0)".

first off, recall that "\displaystyle
f(0,0)=1".

Now, to evaluate "\displaystyle
\lim_{(x,y)\rightarrow(0,0)}f(x,y)", consider the limit along the power curve "\displaystyle
x=t",

"\displaystyle
y=at^2" as "\displaystyle
t\rightarrow 0^+".

Then,

"\displaystyle
\lim_{(x,y)\rightarrow(0,0)}f(x,y)=\lim_{t\rightarrow 0^+}f(t,at^2)=\begin{cases}0 &\text{if }a\in(1,2)\\1&\text{if }a\in(-\infty,1]\cup[2,\infty)\end{cases}"


Since "\lim_{(x,y)\rightarrow(0,0)}f(x,y)=\lim_{t\rightarrow 0^+}f(t,at^2)=\begin{cases}0 &\text{if }a\in(1,2)\\1&\text{if }a\in(-\infty,1]\cup[2,\infty)\end{cases}"

depends on a, thus we have that "\displaystyle
\lim_{(x,y)\rightarrow(0,0)}f(x,y)" does not exist.

"\Rightarrow f(x,y)" is not continuous at "\displaystyle
(0,0)."

"\Rightarrow f(x,y)" is not differentiable at "\displaystyle
(0,0)."